OpenStudy (anonymous):

1/9(36+45q)=2/3(5q+31)

OpenStudy (anonymous):

3(5q+31)=18(36+45q) 5q+31=6(36+45q) 5q+31=216+270q -185=265q q=-185/265 q=37/53

OpenStudy (anonymous):

actually, its -37/53